9. Household income does not tend to follow a normal distribution in a particular state, yet average income is approximately $45,000/year in this state, with a standard deviation of about $9000. Within what income range would approximately 93.75% of the household incomes lie within this state? Show all work as to how this is obtained.
Can someone please help me with this, I don't know what to do at all. I need some explanation please

Answer by stanbon(57377)

Household income does not tend to follow a normal distribution in a particular state, yet average income is approximately $45,000/year in this state, with a standard deviation of about $9000.
-----------
Within what income range would approximately 93.75% of the household incomes lie within this state? Show all work as to how this is obtained.
-------------------------------------------------
Draw the picture of a normal curve with 93.75% centered on the mean.
---
The two tails = 1- 0.9374 = 0.0626
---
The left tail is 0.0626/2 = 0.0313
----
Find the z-value with a left tail of 0.0313
z = invNorm(0.0313) = -1.8620
---
Find the corresponding income value:
x = zs + u
x = -1.8620*9000 + 45000 = $28,242.00
---
Similarly the upper income level is
x = +1.8620*9000 + 45000 = $61,758
======================================
Cheers,
Stan H.
===========